OpenStudy (anonymous):

What is the solution of the equation? 2 5^square root (x+6)^3+3=19

OpenStudy (anonymous):

not so easy to read this one \[25\sqrt{x+6}^3+3=19\]?

OpenStudy (anonymous):

\[2 \sqrt[5]{(x+6)^{3}} +3 =19\]

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

fifth root right?

OpenStudy (anonymous):

2^5 sqrt((x+6)^3) + 3 = 19 32 sqrt ((x+6)^3) = 16 sqrt ((x+6)^3) = 0.5 (x+6)^3 = 0.25 x+6 = 0.6299605249 x = -5.3700394751 The answer is nothing.

OpenStudy (misty1212):

first subtract 3 from both sides

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

what @misty1212 said, subtract 3 as a first step what do you get?

OpenStudy (anonymous):

Answer choices are A.26 C.14 D.2

OpenStudy (anonymous):

A

OpenStudy (anonymous):

yeah yeah we can get the answer subtract 3 and get \[2\sqrt[5]{x+6}=16\] then divide both sides by \(2\)

OpenStudy (anonymous):

Subtract from 19 or from the whole equation itself?

OpenStudy (anonymous):

from both sides your goal is to get \(\sqrt[5]{x+6}\) by itself

OpenStudy (anonymous):

16/2=8

OpenStudy (anonymous):

just like if you were solving \[2x+3=19\] for \(x\) you would a) subtract 3 then b) divide by 2

OpenStudy (anonymous):

idk how to divide the other side

OpenStudy (anonymous):

right

OpenStudy (anonymous):

you have it \[\sqrt[5]{x+6}^3=8\]

OpenStudy (anonymous):

now to get \(x\)

OpenStudy (anonymous):

I just dont know what to do with the left side lol

OpenStudy (anonymous):

I appreciate you taking your time to help me btw

OpenStudy (anonymous):

lets start with \[\sqrt[5]{(x+6})^3=8\]

OpenStudy (anonymous):

getting there was the first two steps, you good til here (yw btw)

OpenStudy (anonymous):

\[\sqrt[5]{(x+6})^{\color{red}3}=8\] lets get rid of that \(\color{red}3\) next it is an exponent, so to get rid of it, take the cubed root of both sides you know what the cubed root of 8 is?

OpenStudy (anonymous):

I have no clue

OpenStudy (anonymous):

think of a number that, when you multiply it by itself 3 times, you get 8 just guess, you have very few choices if you get it wrong, i will tell you

OpenStudy (anonymous):

2

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

so taking the cubed root of both sides gets rid of that exponent of 3 on the left and \[\sqrt[5]{(x+6)^3}=8\] becomes \[\sqrt[5]{x+6}=2\]

OpenStudy (anonymous):

now we have to get rid of the index of 5, which we do by raising both sides to the power of 5 what is \(2^5\)?

OpenStudy (anonymous):

you can back solve it

OpenStudy (anonymous):

i didn't lose you there did it?

OpenStudy (anonymous):

Sorry catching up my computer froze

OpenStudy (anonymous):

yeah it does that

OpenStudy (anonymous):

os i mean

OpenStudy (anonymous):

just put all the options one by one in the question and one option will satisfy the equation

OpenStudy (anonymous):

raisiing both sides to the power of 5 gives \[x+6=2^5\] or \[x+6=32\]

OpenStudy (anonymous):

32

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now subtract 6 from both sides and you are done

OpenStudy (anonymous):

32-6

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

26

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Wow thank you, I would have never figured that out without you breaking it down

OpenStudy (anonymous):

you are quite welcome

OpenStudy (anonymous):

I gave you a medal, well deserved

OpenStudy (anonymous):

I was right the answer is 26

OpenStudy (anonymous):

You were right, but you just gave the answer instead of explaining lol that wont help me on a test

OpenStudy (anonymous):

i prefer quick learning

OpenStudy (anonymous):

as like in mcq's u have to save time....not to write papers for one question only